Isaac Inc. began operations in January 2016. For certain of its property sales, Isaac recognizes income in the period of sale for financial reporting purposes. However, for income tax purposes, Isaac recognizes income when it collects cash from the buyer's installment payments.

In 2016, Isaac had $600 million in sales of this type. Scheduled collections for these sales are as follows:

2016 $60 million
2017 120 million
2018 120 million
2019 150 million
2020     150 million
$600 million


Assume that Isaac has a 30% income tax rate and that there were no other differences in income for financial statement and tax purposes.

Ignoring operating expenses and additional sales in 2017, what deferred tax liability would Isaac report in its year-end 2017 balance sheet?

A-$54 million.

B-$144 million.

C-$126 million.

D-$180 million.

For its first year of operations, Tringali Corporation's reconciliation of pretax accounting income to taxable income is as follows:

Pretax accounting income $300,000
Permanent difference      (15,000)
285,000
Temporary difference-depreciation      (20,000)
Taxable income    $265,000


Tringali's tax rate is 40%. Assume that no estimated taxes have been paid.

What should Tringali report as its deferred income tax liability as of the end of its first year of operations?

A-$35,000.

B-$20,000.

C-$14,000.

D-$8,000.

Pretax accounting income for the year ended December 31, 2016, was $50 million for Truffles Company. Truffles' taxable income was $60 million. This was a result of differences between straight-line depreciation for financial reporting purposes and MACRS for tax purposes. The enacted tax rate is 30% for 2016 and 40% thereafter. What amount should Truffles report as the current portion of income tax expense for 2016?

A-$15 million.

B-$18 million.

C-$20 million.

D-$24 million
Answer #1

Answer: (Amount in $s)

1)

Year 2016

Timing difference resulting into liabilty for the year 2016 = 600 - 60 = 540

Therefore Deferred tax liabilty = 540 * 0.30 = 162

Year 2017

Brought forward Timing difference resulting into liabilty = 540

Timing difference resulting into asset for the year 2017 = 120

Therefore balance timig difference resulting into liability = 540 - 120 = 420

Thus Deferred tax liability for the year 2017 = 420 * 0.30 = 126 million

Answer = (c)

2)

Deferred tax liability as at the end of first year = 20,000 * 0.40 = $8,000.

Answer = (d)

3)

Income tax expense for the tear 2016 = 60 * 0.30 = $18 million.

Answer = (b)